### v1

# Harnessing the Unseen: The Hidden Influence of Intrinsic Knowledge in Long-Context Language Models

Abstract
Recent advances in long-context models (LCMs), designed to handle extremely long input contexts, primarily focus on utilizing external contextual information, often leaving the influence of large language models’ intrinsic knowledge underexplored. In this work, we investigate how this intrinsic knowledge affects content generation and demonstrate that its impact becomes increasingly pronounced as context length extends. Furthermore, we show that the model’s ability to utilize intrinsic knowledge, which we call intrinsic retrieval ability, does not improve simultaneously with its ability to leverage contextual knowledge through extrinsic retrieval ability. Moreover, better extrinsic retrieval can interfere with the model’s ability to use its own knowledge effectively, limiting its full potential. To bridge this gap, we design a simple yet effective Hybrid Needle-in-a-Haystack test that evaluates models based on their capabilities across both retrieval abilities, rather than solely emphasizing extrinsic retrieval ability. Our experimental results reveal that Qwen-2.5 models significantly outperform Llama-3.1 models, demonstrating superior intrinsic retrieval ability. Moreover, even the more powerful Llama-3.1-70B-Instruct model fails to exhibit better performance under LCM conditions, highlighting the importance of evaluating models from a dual-retrieval perspective.
1 Introduction
Recent advancements in open-source models such as LLaMA (Grattafiori et al., 2024), Qwen (Yang et al., 2024a; b) and closed-source models such as GPT-4 (Achiam et al., 2023; OpenAI, 2024), Claude (Anthropic, 2024) and Gemini (DeepMind, 2024) have all incorporated long-context capabilities, extending their context windows to tens of thousands of tokens. For instance, GPT-4o and Claude 3.5 have expanded their context windows to 128K and 200K tokens respectively. The emergence of those Long Context Models (LCMs) has significantly advanced the ability of language models to process and generate coherent content over extended inputs. These models have become increasingly useful for various tasks, including document summarization (Chang et al., 2023; Yen et al., 2025), multi-turn conversations (Maharana et al., 2024) and question answering (Yang et al., 2018; An et al., 2023; Karpinska et al., 2024), where retaining and utilizing long-term dependencies is essential.
Numerous benchmarks have been proposed to evaluate LCMs’ effectiveness in handling long contexts. For example, the Needle-in-a-Haystack test (Kamradt, 2023) inserts critical information (i.e. “needle”) in long irrelevant documents (i.e. “haystack”) and test LCMs’ ability to retrieve the information. Building upon this experiment, benchmarks such as RULER (Hsieh et al., 2024), InfiniteBench (Zhang et al., 2024) and HELMET (Yen et al., 2025) extend to more realistic tasks such as QA, summarization etc., and very long contexts (100K+ tokens). These benchmarks systematically assess LCMs’ capacity to process and generate meaningful outputs over extraordinarily long contexts.
However, the aforementioned benchmarks for LCMs primarily evaluate how well these models leverage external long-context inputs, without considering the influence of their parametric (intrinsic) knowledge (Roberts et al., 2020; Wang et al., 2021; Jiang et al., 2020). Parametric knowledge has been shown to be critical in language models for short context tasks (Xie et al., 2023; Xu et al., 2024b) and retrieval-augmented generation (Mallen et al., 2023; Wang et al., 2024a). In this work, we argue that intrinsic knowledge plays a crucial role in long-context generation and can significantly impact model performance. Concretely, we create a new dataset - IWhoQA - based on WhoQA (Pham et al., 2024) and test the behavior of LCMs when their intrinsic knowledge either supports or conflicts with the extrinic knowledge in the long context input. We show that LCMs consistently achieve better performance when the context supports their intrinsic knowledge and they gradually rely more on intrinsic knowledge when the input context length is increased.
Given the importance of intrinsic knowledge for LCMs, we further seek to determine whether enhancing a model’s extrinsic retrieval ability can influence the expression of intrinsic knowledge. Specifically, we compare between RoPE (Su et al., 2021) and STRING (An et al., 2024b), a modified version of RoPE designed for long-context tasks, on a curated dataset collected from HotPotQA (Yang et al., 2018). We show that STRING improves extrinsic retrieval ability but unintentionally suppresses intrinsic retrieval ability, even when the latter positively contributes to generation quality. This finding highlights a trade-off between utilizing external long-context inputs and leveraging the LCMs’ intrinsic knowledge.
Based on these observations, we propose a simple yet efficient Hybrid Needle-in-a-Haystack test to offer a comprehensive assessment of how well models can integrate the intrinsic and extrinsic retrieval ability during generation. Specifically, we design a set of questions such as “What’s the favorite thing of the person who wrote {Book_Name}?” which require the model to first retrieve the author’s name based on intrinsic knowledge and then retrieve a needle to answer the question. Our experimental results show that the ability of the Qwen2.5 family to integrate both intrinsic and extrinsic knowledge scales linearly with model size. In contrast, the Llama3.1 family fails to achieve noticeable improvement even with a substantial increase in model parameters, indicating that Llama3.1 struggles to effectively leverage its intrinsic knowledge in long-context scenarios. Our contributions can be summarized as follows:
  • We present the first study demonstrating that the intrinsic knowledge of LCMs plays a significant role in long-context generation, with its influence becoming increasingly pronounced as the context length grows.
  • We show that naively improving extrinsic retrieval ability will unintentionally suppress intrinsic retrieval ability, revealing a trade-off between utilizing external inputs and leveraging intrinsic knowledge in LCMs.
  • We introduce a novel Hybrid Needle-in-a-Haystack test to examine LCMs’ ability to integrate the intrinsic and extrinsic retrieval ability during generation. We show that this ability doesn’t always scale with the model size, highlighting limitations in existing LCMs.
2 Related Work
Long Context Models
Since training transformers on long context from scratch is prohibitively expensive, research has focused on extending the context window of existing transformers through two broad categories: 1) Finetuning on a small amount of long context data (Peng et al., 2024; Lazarevich et al., 2025; Chen et al., 2023; jquesnelle, 2023) and 2) Training-free methods based on reusing relative positions seen during training (An et al., 2024a; bloc97, 2023; emozilla, 2023; Su, 2023; Jin et al., 2024). Self-Extend (Jin et al., 2024), Dual Chunk Attention (An et al., 2024a), and STRING (An et al., 2024b) all propose replacing large relative distances between query and key tokens by overwriting them with proportionately smaller relative distances seen during training. Relative distances below a certain threshold remain unmodified.
Another challenge for long context language modeling is storing key and value activations for all past tokens (the KV cache). Significant research has focused on selectively attending to only parts of the context with sparse attention (Han et al., 2024; Xiao et al., 2024; Jiang et al., 2024). Additional optimization comes from estimating the importance of the KV cache and evicting low-importance cache entries (Fu et al., 2024; Feng et al., 2025; Cai et al., 2024; Li et al., 2024).
Long Context Evaluation
Multiple synthetic tasks have been proposed to evaluate LCMs, including Needle-in-a-Haystack (NIAH) test (Kamradt, 2023), long-context QA, summarization, reasoning, extrinsic learning (ICL) and retrieval-augmented generation (RAG), among others (Hsieh et al., 2024; Kim et al., 2025; Gao et al., 2025; Liu et al., 2024; Yen et al., 2025). Variants of NIAH include multiple-needle, long-needle, and needle-in-needle configurations (Gao et al., 2025) and have also been extended to test long-context reasoning (Kuratov et al., 2024), mathematical (Wang et al., 2024b) and multilingual capabilities (Hengle et al., 2024). However, several works have found that the standard NIAH task may not truly reflect an LCM’s downstream capabilities and have proposed comprehensive benchmarks such as RULER (Hsieh et al., 2024) and it’s multilingual extension ONERULER (Kim et al., 2025), LongBench (Bai et al., 2024) , LongGenBench (Liu et al., 2024), L-Eval (An et al., 2023), InfiniteBench (Zhang et al., 2024) and HELMET (Yen et al., 2025). These benchmarks contain several tasks such as citation generation, reranking, QA, summarization, ICL, and RAG. Karpinska et al. (2024) introduced NoCha, a dataset of human-written true and false claims about recently published fictional books. While previous work has primarily studied long context in isolation, some going so far as to use entirely fictional contexts (Gao et al., 2025; Karpinska et al., 2024), to the best of our knowledge the problem of how an LLM’s intrinsic knowledge affects its long context abilities has not been explored.
Intrinsic Knowledge
During pre-training, LLMs memorize a vast amount of knowledge into their parameters, known as parametric (intrinsic) knowledge (Roberts et al., 2020; Wang et al., 2021; Jiang et al., 2020). This parametric knowledge can often conflict with new information presented via prompting through a phenomenon known as knowledge conflict (Xie et al., 2023; Longpre et al., 2022; Pan et al., 2023; Xu et al., 2024a). While earlier works have found that LLMs rely on parametric knowledge when prompted with simple counter-factual statements (Longpre et al., 2022; Chen et al., 2022), more recent research have demonstrated that LLMs rely more on the information presented as the context (Xie et al., 2023; Pan et al., 2023; Wan et al., 2024). Pan et al. (2023) uncovers LLMs’ sensitivity to small amounts of even misinformation in context and Xu et al. (2024a) applies prompting to persuade LLMs to disregard their parametric knowledge. However, Xie et al. (2023) further finds that when presented with conflicted evidence, LLMs resolve in favor of parametric knowledge. Zhou et al. (2023) proposes framing the conflicting information as user opinion to prevent the LLM from taking it as fact. We refer the reader to Xu et al. (2024b) for further details. However, to the best of our knowledge, the relation between parametric knowledge and long-context tasks has not been studied.
3 The Role of Intrinsic Knowledge in Long-Context Generation
Current benchmarks for LCMs focus primarily on assessing a model’s ability to extract important information or perform reasoning based only on the provided context. In contrast, we seek to investigate the role of the LCMs’ intrinsic knowledge. This section begins with the question: What role does the intrinsic knowledge of LCMs play in long-context generation? We aim to explore how the model’s intrinsic knowledge interacts with contextual data, potentially influencing the quality and accuracy of the generated results.
3.1 Dataset Creation
We design new dataset to first investigate whether LCMs naturally incorporates intrinsic knowledge during generation. Recognizing that existing datasets within the knowledge-conflict (Xie et al., 2023; Xu et al., 2024b; Pham et al., 2024) also aim to identify and investigate intrinsic knowledge of LLMs, we chose WhoQA (Pham et al., 2024), a QA dataset which contains several context-answer pairs for each question, as our initial dataset.
For each example in WhoQA, we generated multiple answers based on different phrasings of question and retained only those examples where the model generates only one stable answer. This methodology ensures that the intrinsic knowledge in the model contains only one answer for this example. Given that the intrinsic knowledge may vary among different LCMs, we collect separate datasets for each LCMs. Consequently, we constructed datasets comprising 300 examples for each model, which we refer to as the I-WhoQA dataset.
As our objective is to assess how intrinsic knowledge affects long-context generation, we select two different context-answer pairs for each example in I-WhoQA: (1) Context that aligns with the model’s intrinsic knowledge, referred to as I-WhoQA-internal subset. (2) Context that conflicts with the model’s intrinsic knowledge, referred to as I-WhoQA-conflict subset. By using these two different subsets from I-WhoQA dataset, we can uncover how intrinsic knowledge influences the generation in long-context scenarios. See Appendix 6 for an example of our I-WhoQA dataset.
Figure 1: Left) LCMs struggle to retrieve the answer from the context on the I-WhoQA-Conflict-subset i.e. when the context information conflicts with its intrinsic knowledge. Right) The upward trend of I-WhoQA-conflict (red) shows that when the intrinsic knowledge conflicts with the context, the likelihood of a LCM relying on intrinsic knowledge steadily increases with larger contexts.
3.2 Experiment
Experimental Setting
To validate the role of intrinsic knowledge in long-context generation, we conducted analyses using two curated I-WhoQA subsets mentioned above, alongside two different backend models — mistral-7b-instruct-v0.2 (Jiang et al., 2023) and Llama3.1-8B-instruct (Grattafiori et al., 2024). For each backend model, we used greedy decoding strategy with a maximum of 32 generation tokens. To examine the impact of context lengths, which is key factor in long-context generation, we report experimental results with varying context lengths. For the same example, the only difference across these different context settings is the amount of irrelevant contents inserted into the context, which is consistent with the setup of Needle-in-a-Haystack test.
Main Results
The results in Figure 1 show that the intrinsic knowledge of LCMs will influence its generation and can be more significant with the increase of context length. Firstly, the left side of Figure 1 (Answer Matches Context) shows the performance on the IWhoQA-conflict is consistently worse than on IWhoQA-internal subset. This demonstrates that the LCMs are struggling to answer according to the context when its intrinsic knowledge contains a different answer. Secondly, from the right side of Figure 1 (Answer Conflict With Context) on the I-WhoQA-conflict subset, where conflicting contexts are provided as supporting facts, we observed a notable increase in the proportion of outputs adhering to intrinsic knowledge as context length increased. This trend is especially pronounced when Llama-3.1-8B-Instruct serve as the backend model, which can handle longer context lengths.
4 The Trade-off Between Intrinsic and Extrinsic Retrieval Abilities
In the previous section, we demonstrated that the intrinsic knowledge within LCMs plays a crucial role in long-context generation. This influence becomes increasingly significant as the context length grows. While existing research (An et al., 2024b; Han et al., 2024; emozilla, 2023) on long-context modeling has largely focused on enhancing models’ extrinsic retrieval abilities, there has been little analysis of the influence of LCMs’ intrinsic knowledge during long-context generation, which we refers to as intrinsic retrieval ability. In this section, we aims to solve the question: Will the expression of intrinsic knowledge be influenced by the extrinsic retrieval ability in long-context generation? We study the effect of STRING (An et al., 2024b), an improved position encoding method, on extrinsic and intrinsic retrieval ability of LCMs. Our experimental results reveal that simply enhancing extrinsic retrieval ability with STRING will inadvertently alter the expression of intrinsic knowledge, which in some cases can lead to a decline in model performance.
4.1 Dataset Creation
In Section 3, we built the I-WhoQA dataset, ensuring that all examples can be answered using the model’s intrinsic knowledge. Based on the I-WhoQA dataset, we further constructed the I-WhoQA-irrelevant subset. In this subset, instead of providing context relevant to the current question, we inserted completely irrelevant context to assess how models handle unrelated information and if STRING can handle the situation that need intrinsic retrieval ability rather than extrinsic retrieval ability.
The I-WhoQA-irrelevant subset represents a relatively extreme setting and may occur infrequently in real-world scenarios. More commonly, the context may contain content that is related to the question but cannot actually be used to derive the correct answer. To address this issue, we construct a corresponding dataset from HotpotQA dataset (Yang et al., 2018), where the context includes seemingly relevant but ultimately unhelpful information for answering the question. Based on the HotpotQA dataset, we construct two subsets by examining the relationship between the paragraphs provided by HotpotQA, the model’s intrinsic knowledge and the reference answer.
First, to distinguish the role of provided context from that of the model’s intrinsic knowledge, we ensured that the answers derived from the given paragraphs and the model’s intrinsic knowledge are not the same. Building on this distinction, we obtained two final subsets. The first subset, referred to as the HotpotQA-context subset, contains examples where the reference answer can be derived from the given passages rather than the intrinsic knowledge. The second subset, known as the HotpotQA-internal subset, consists of examples where the reference answer matches the model’s intrinsic knowledge. We constructed a dataset consisting of 200 examples for each subset, except for the HotpotQA-internal subset for Llama3.1-8B-Instruct, which contains a total of 93 examples.
Figure 2: I-WhoQA-irrelevant subsets: LCMs are often unable to ignore context not relevant to the question and recall the answer from their intrinsic knowledge. As the length of the context is increased, LCMs gradually start ignoring the context and generating according to their intrinsic knowledge. However, using enhanced positional encoding designed for long context tasks such as STRING causes LCMs to re-focus on irrelevant context and generate wrong answers.
Figure 3: Performance of LCMs on HotPotQA-context and -internal subsets. Left) STRING improves performance on the HotPotQA-context subset by enhancing extrinsic retrieval ability. Right) STRING hinders the model’s ability to recall intrinsic knowledge and leads to the decrease of performance.
4.2 Experiment
Experimental Setting
We use the same backend models and generation configuration as described in Section 3, applied to the newly constructed datasets. For the STRING method, we configure the local value to 128 and two different the shift-ratios: 0.25 and 0.33. We report the average performance across these two settings. More details about STRING can be found in An et al. (2024b).
Main Results
Figure 2 shows that STRING consistently achieves lower exact match accuracy than the baseline RoPE method. We hypothesize that this degradation may result from STRING enhancing the extrinsic retrieval ability under long-context scenarios. As a result, even when the provided context is irrelevant, the STRING method will still attempts to retrieve and rely information from the context rather than using the model’s intrinsic knowledge that contains the correct answer.
Results on HotpotQA-context subset and HotpotQA-internal subset are shown in Figure 3. STRING significantly improves performance on the HotpotQA-context subset, where the model needs to rely on the provided context to derive the reference answer. The improved performance implies that STRING is improving extrinsic retrieval ability on LCMs which supports our hypothesis above. However, on the HotpotQA-internal subset where the context is related but not useful, we find STRING underperforms RoPE which matches our finding on I-WhoQA-irrelevant. In conclusion, the enhanced extrinsic retrieval ability of STRING interferes with the expression of the model’s intrinsic knowledge, resulting in worse performance compared to the vanilla models. This trend becomes even more apparent on longer context length since STRING’s boosted extrinsic retrieval ability can be more prominent with longer contexts as shown on the Llama-3.1-8B-Instruct HotPotQA-context subset.
5 Hybrid Needle-in-a-Haystack Framework
So far, we have confirmed that intrinsic knowledge plays a crucial role in the generation process of LCMs, but the intrinsic retrieval ability does not improve in parallel with the extrinsic retrieval ability. In this section, we want to examine LCMs’ ability to integrate the intrinsic and extrinsic retrieval ability during long-context generation. Previous benchmarks (Hsieh et al., 2024; An et al., 2023; Zhang et al., 2024) have primarily focused on the effective utilization of external context and overlooked assessing the impact of LCMs’ intrinsic knowledge. To address this gap, we designed a simple yet effective Hybrid Needle-in-a-Haystack (NIAH) test aimed at simultaneously testing intrinsic and extrinsic retrieval ability, providing a more comprehensive assessment of how these two capabilities interact, complement, or potentially interfere with one another, ultimately affecting the quality of generation.
5.1 Dataset Creation
NIAH test uses irrelevant content as a haystack and inserts the answer from a question-answer pair into the haystack to evaluate LCMs’ extrinsic retrieval ability under different context lengths and insertion depths. Unlike the original NIAH, our Hybrid NIAH test simultaneously tests both retrieval abilities by integrating scenarios where the model needs to first utilize its intrinsic knowledge and then perform extrinsic retrieval on the haystack. A comparison between NIAH and Hybrid NIAH is illustrated in Figure 4. The Hybrid NIAH introduces an extra step designed to trigger the intrinsic retrieval ability, requiring the model to recall the relevant target from intrinsic knowledge. The model must use this recalled target to perform extrinsic retrieval from the provided context, ultimately producing the final output.
Firstly, to implement the intrinsic retrieval step, we use the I-WhoQA dataset constructed in Section 3. To ensure that the question-answer pairs only require knowledge that exists across all models, we perform an intersection operation between the I-WhoQA datasets obtained from Mistral-7B-Instruct-v0.2 and Llama3.1-8B-Instruct. Secondly, for the extrinsic retrieval step, we adopt a similar approach to the NIAH test and create various random facts to serve as inserted answers to the corresponding questions. For the remaining irrelevant haystack construction, we also follow the NIAH test and use the PaulGrahamEssays dataset  [^1] https://huggingface.co/datasets/sgoel9/paul_graham_essays to construct the target length haystack.
To prevent the model from exploiting potential syntactic patterns present in the question-answer pairs to extract the inserted “needle” from the context, we introduce varying amounts of random facts as interference (from 0 to 3). This additional noise ensures that models understand the task and the context rather than relying on superficial patterns, thereby enhancing the robustness and reliability of the final result. Details of the final knowledge pairs are listed in Appendix A.
Figure 4: Upper) Needle-in-a-Haystack. It involves directly inserting the answer into the haystack and retrieving it. Lower) Hybrid Needle-in-a-Haystack. It requires a two-step process: first, performing intrinsic retrieval to identify the retrieval target based on the model’s intrinsic knowledge, and then retrieving the answer from the haystack.
5.2 Experiment
Experimental Setting
We evaluated models from three different families with various specifications: Mistral, Llama3.1, and Qwen2.5. For all models, we utilized the Instruct versions and employed greedy decoding strategy. We set the maximum generation length to 32 and 64. For each model, we divided the maximum context length into 40 intervals. For each interval, we defined 10 progressively increasing insertion depths and conducted Hybrid NIAH tests accordingly. This setup allows us to systematically evaluate the models’ performance across varying input lengths and insertion depths. To ensure the stability of the results, we use multiple examples and report the average result. Finally, to measure the similarity between the predictions and the references, we adopt the method provided by PyramidKV (Cai et al., 2024) [^2] https://github.com/Zefan-Cai/KVCache-Factory/, which calculates the final result based on exact matches between tokens. The formula is as follows:
$Score=\frac{|P|\cap|A|}{|A|}$
where $P$ and $A$ are tokens from predictions and references respectively. More detail settings are listed in Appendix Table 2.
Model Generation Length = 32 Generation Length = 64
Random Facts Random Facts
0 1 2 3 0 1 2 3
Needle-in-a-Haystack
Mistral-7B-Instruct-v0.2 100 100 99.55 98.41 - - - -
Llama-3.1-8B-Instruct 100 99.21 99.33 98.83 - - - -
Qwen2.5-7B-Instruct 99.78 99.64 99.81 99.92 - - - -
Hybrid Needle-in-a-Haystack
Mistral-7B-Instruct-v0.2 89.31 58.30 53.53 53.21 97.83 63.28 58.16 58.11
Mistral-7B-Instruct-v0.3 73.17 59.54 56.97 56.94 76.56 62.48 58.01 57.73
Llama-3.1-8B-Instruct 92.97 83.66 72.60 67.47 96.71 90.28 80.93 74.39
Llama-3.1-70B-Instruct 95.73 86.39 74.48 71.81 95.77 89.88 77.66 74.68
Qwen2.5-7B-Instruct 86.38 83.32 77.23 73.14 90.72 88.46 80.98 77.03
Qwen2.5-14B-Instruct 54.46 76.74 75.32 73.97 92.95 94.32 89.85 85.60
Qwen2.5-72B-Instruct 64.01 98.13 97.92 97.77 94.16 99.53 99.84 98.86
Qwen2.5-7B-Instruct-1M-128k 97.72 84.83 82.10 79.31 98.50 84.89 82.22 79.37
Qwen2.5-7B-Instruct-1M 60.07 48.09 43.70 42.91 60.77 49.26 43.43 43.21
Table 1: Average accuracy score of different models. Qwen2.5-7B-Instruct-1M-128k means we set the max input length to 128k.
Main Results
The averaged results for the Hybrid NIAH test are presented in Table 1 and we list the detailed results in Appendix B. From these results, we can observe three interesting findings. Firstly, our Hybrid NIAH test is significantly more challenging than the standard NIAH test, especially on the long-context scenarios as shown in the results on Qwen2.5-7B-Instruct-1M in Figure 5. Moreover, by inserting varying amounts of random facts, we observe a gradual decline in model performance, while the NIAH test results remain unaffected by this interference as shown in Table 1. This discrepancy indicates that models tend to rely on superficial patterns within the context to extract answers when there is a single needle, rather than genuinely leveraging their intrinsic knowledge. The insertion of random facts effectively disrupts this pattern-based retrieval process, allowing us to better assess whether a model truly utilizes its intrinsic knowledge rather than merely exploiting contextual shortcuts. Multiple-needle Hybrid NIAH test serves as a more rigorous evaluation of how well models integrate their intrinsic retrieval ability with the long-context generation.
Secondly, models from the Mistral and Llama3.1 families fail to achieve significant improvements on the Hybrid-NIAH test, even when upgraded to larger or more advanced versions. This suggests that these models struggle to effectively utilize their intrinsic knowledge under long-context scenarios. Despite increased model capacity, their performance remains constrained when external context is incomplete or ambiguous. In contrast, the Qwen2.5 family shows substantial improvements as the model scales up. Larger Qwen2.5 models are more capable of effectively leveraging their intrinsic knowledge, providing more accurate and factually consistent results even when the external context is incomplete.
Figure 5: Hybrid NIAH test results. Upper) Qwen2.5-7B-Instruct-1M with generation length 32. Lower) Qwen2.5-72B-Instruct with generation lengths 32 and 64. The number of random facts was set to 0.
Finally, we observe an interesting phenomenon in the Qwen2.5 models: when the generation length is set to 32, both Qwen2.5-14B-Instruct and Qwen2.5-72B-Instruct perform worse that Qwen2.5-7B-Instruct on the single-needle Hybrid NIAH test. However, when the generation length is increased to 64, their performance improves dramatically as shown in Figure 5. Upon inspecting the outputs manually, we find that these models often first explicitly state their inability to retrieve the answer from the context before producing the correct result. Interestingly, when we introduce several random facts, the models can directly answer the corresponding question without hesitation as shown in Appendix Figure 7. We hypothesize that this behavior is closely related to the training process of Qwen2.5 family. Firstly, when only a single fact is present, the model’s chunked attention mechanism (An et al., 2024a) struggle to effectively focus on the inserted needle when generating the first token. As a result, larger Qwen variants initially rely on their logical reasoning to realize that they do not know the answer and start generating a refusal. But as each refusal token is generated, the model can attend to the context again to find the needle and generate the correct answer even through it refused initially. In contrast, the smaller Qwen-2.5-7B-Instruct model starts hallucinating an answer although it’s unsure instead of outright refusing. Secondly, when multiple needles are inserted, the syntactical similarity between the needles help the larger Qwen models to focus on the context and they do not refuse answering initially. This observation suggests that the Qwen2.5 attention mechanism is highly sensitive to pattern similarity, and syntactical structure can help guide the model’s attention to the relevant part from the context. We leave further investigation of this phenomenon for future work.
6 Conclusion
In this paper, we reveals that the intrinsic knowledge within LCMs plays a crucial role in generation, particularly under long-context scenarios. We demonstrate that enhancing extrinsic retrieval ability can inadvertently suppress the model’s ability to utilize its intrinsic knowledge, highlighting a fundamental trade-off between these two retrieval mechanisms. We further propose a Hybrid Needle-in-a-Haystack test, providing a comprehensive framework to evaluate both intrinsic and extrinsic retrieval ability. Our findings show that even large LCMs struggle to combine the two retrieval abilities and cannot fully leverage its intrinsic knowledge during the long-context generation.
Supplementary Material: Appendices
Figure 6: One example on I-WhoQA dataset on Mistral-7B-Instruct-v0.2.
Model Num of Examples Max Context Length Num of Depths Num of Intervals
Mistral-7B-Instruct-v0.2 5 32768 40 10
Mistral-7B-Instruct-v0.3 2 32768 40 10
Llama-3.1-8B-Instruct 5 131072 40 10
Llama-3.1-70B-Instruct 2 131072 40 10
Qwen2.5-7B-Instruct 5 131072 40 10
Qwen2.5-14B-Instruct 2 131072 40 10
Qwen2.5-72B-Instruct 2 131072 40 10
Qwen2.5-7B-Instruct-1M-128k 5 131072 40 10
Qwen2.5-7B-Instruct-1M 2 1010000 40 10
Table 2: NIAH and Hybrid NIAH test configurations for different backend models.
Figure 7: Hybrid NIAH examples from Qwen2.5-72B-Instruct. The generation length was set to 64.
Appendix A Hybrid Needle-in-a-Haystack Knowledge Dataset
In the WhoQA dataset, various entity types are provided, including [date of birth], [author], [country], [date of death], [father], [composer], [performer], [genre], [creator], [occupation], [spouse], [publisher], and [mother]. For constructing the question-answer pairs used in Hybrid NIAH test, we focus specifically on the [author] entity type as the basis for our evaluations.
All entities used in Hybrid NIAH tests are illustrated in Figure 3, with the bolded entries representing the target entities used for generating the desired answers. To construct the random facts required for our Hybrid NIAH tests, we randomly select entities from all entities other than the current target entity.
Book Author Book Author
The Stranger Albert Camus The Social Contract Jean-Jacques Rousseau
Antigone Sophocles Crime and Punishment Fyodor Dostoevsky
Leviathan Thomas Hobbes Iphigenia in Tauris Euripides
Dracula Bram Stoker Dover Beach Matthew Arnold
The Birds of Australia John Gould Travels with My Aunt Graham Greene
Conan the Barbarian Robert E. Howard Candide Voltaire
Politics Aristotle Time’s Arrow Martin Amis
What’s the Matter with Kansas? Thomas Frank Bad Science Ben Goldacre
On Liberty John Stuart Mill A Clash of Kings George R. R. Martin
Tipping the Velvet Sarah Waters Shakuntala Kalidasa
The Tale of Genji Murasaki Shikibu The Age of Reason Thomas Paine
The Art of War Sun Tzu Ulysses James Joyce
The Prince Niccolò Machiavelli Lady Lazarus Sylvia Plath
Carmen Georges Bizet The Age of Innocence Edith Wharton
The Temple of Elemental Evil Gary Gygax James Bond Ian Fleming
Histories Herodotus The Fishermen Chigozie Obioma
Pygmalion George Bernard Shaw The Three Musketeers Alexandre Dumas
The Hiding Place Corrie ten Boom The Wonderful Wizard of Oz L. Frank Baum
Misery Stephen King Richard III William Shakespeare
The Nine Billion Names of God Arthur C. Clarke Amazing Grace John Newton
The Seagull Anton Chekhov Paradise Lost John Milton
There Will Come Soft Rains Ray Bradbury The Betrothed Alessandro Manzoni
Parable of the Sower Octavia Butler The Hollow Men T.S. Eliot
Table 3: Knowledge pairs selected from I-WhoQA dataset.
Appendix B Hybrid Needle-in-a-Haystack Results
We listed Hybrid NIAH test results on different models. Figure 8 shows the test results on models from Mistral family. Figure 9 shows the test results on models from Llama3.1 family. Figure 11 shows the test results on models from Qwen2.5 family except Qwen2.5-7B-Instruct-1M. Figure 11 shows the test results on Qwen2.5-7B-Instruct-1M.
Figure 8: Hybrid Needle-in-a-Haystack test results on models from Mistral family. Upper) Mistral-7B-Instruct-v0.2 and Lower) Mistral-7B-Instruct-v0.3.
Figure 9: Hybrid Needle-in-a-Haystack test results on models from Llama3.1 family. Upper) Llama3.1-8B-Instruct and Lower) Llama3.1-70B-Instruct.
Figure 10: Hybrid Needle-in-a-Haystack test results on models from Qwen2.5 family. Upper) Qwen2.5-7B-Instruct, Middle) Qwen2.5-14B-Instruct Lower) Qwen2.5-70B-Instruct.
Figure 11: Hybrid Needle-in-a-Haystack test results on Qwen2.5-7B-Instruct-1M. Upper) Qwen2.5-7B-Instruct-1M-128k, Lower) Qwen2.5-7B-Instruct-1M.